A solution of 2.50 M was diluted by a factor of 0.06, what is the new concentration of the solution?
Answer #1

DIlution factor = 0.06

So the concentration after dilution = 0.06 * 2.50 M = 0.15 M

In analytical chemistry, dilution factor = final volume / initial volume

However here it is not specified the final and initial volumes. Hence we need to assume the dilution factor is 0.06 times dilution.

Hence, the concentration of the final solution = 0.06* initial concentration = 0.06*2.50M = 0.15 M